Swanmore apiary report – August 2021

Here are the latest reports from Swanmore on the progress of our apiary. As of 9 August, four of our five colonies are between queens following recent swarms and splits. Fingers crossed that all colonies will be queen-right by the end of the month so they can concentrate on their winter preparations!

Hive 5 (George’s)

Lets look at George’s hive (number 5) first, because it’s the start of the story for hive 1 too! Until Monday, the colony was on double brood and had four well-stocked supers.

When opened, multiple queen cells were found, most of them charged. Denise and Richard estimated the larvae inside were four or five days old; so a couple more days and a swarm would have gone. A very fortuitously-timed inspection!

Hive 5 (position 9)

Because of the colony’s strength (not to mention the weight of it on the stand) Denise and Richard decided to split the two brood boxes to create separate colonies. The upper brood box (including the queen) was moved to position 1 in the apiary.

One strong-looking queen cell was left in hive 5 but the bees may now raise further queen cells from eggs. If they do, these can be removed to prevent a secondary (cast) swarm so the hive will be checked again within seven days.

One full super of honey was removed from hive 5 and given to the new hive 1.

Hive 5 Update: 9 August

Richard working hive 5, 9 August

Since the last inspection, one week ago, the bees have been busy creating multiple queen cells to add to one left deliberately in the hive by the beekeepers. Most of these cells were on one frame.

There are three supers on the hive which between them were about 60% full. In the brood box, there was also one frame with capped honey on both sides.

We also found nearly three frames of capped brood but as expected, no unsealed brood or eggs were seen. We decided to use the frame with multiple cells to make-up a three frame nuc and we additionally transferred to the nuc, the frame of capped honey and another containing capped brood.

Finally, we shook in some extra bees for good measure and added another frame of drawn comb. The nuc has now been transferred to the West End apiary.

Back in Hive 5, we removed most of the queen cells on other frames leaving only two. We will now leave the bees alone until at the end of the month when hopefully, they’ll be queen right and can get in with their winter preparations.

Note, hive 5 needs five new frames of foundation which will be added within the next couple of days. This is to replace the frames taken for the nuc and some dirty comb.

Hive 1

Colony created on Monday 2 August. Brood box containing brood, stores and the laying queen removed from hive 5 (George’s) and placed on position 1 in the apiary.

It was noted that this is a very productive queen and she has been earmarked for the breeding programme next year provided we can get her through winter.

One full super was removed from hive 5 and placed on hive 1 to ensure new colony is adequately provisioned.

The box removed from hive 5 contained five frames of sealed brood. There were also seven queen cells (five of which were charged) and these were destroyed to remove the immediate possibility of swarming.

Hive one is made up of a new floor, queen excluder, crown board, new roof (all association equipment) but the brood box belongs to George.

Hive 1, position 1

This colony will be checked again within seven days in case the bees are still feeling swarmy and produce more queen cells.

Hive 1 Update: 9 August

Still plenty of capped brood in the newly created hive 1.

Although the population of Hive 1 wasn’t noticeably reduced (seven days after colony split), it was soon apparent a swarm had gone. The weather was changeable with occasional showers but the bees were perfectly calm throughout allowing for a thorough inspection. We checked through the frames three times but the queen (marked white) had gone. What a shame! There was no uncapped brood or eggs either.

It seems that the bees swarm preparations were too advanced to stop and they continued with their original plan despite last week’s manipulation. Based on the estimated age of the queen cells last week, the swarm probably issued on Tuesday or Wednesday.

In all, four emergency cells were found which we reduced to two. There was a good level of stores both in the brood box and super, although the bees seem to be moving stores from the super to the brood box. We concluded this because there was plenty of very freshly capped honey is the brood box and a reduced amount in the supers.There was also capped brood on five frames.

The hive entrance was set to small because the beekeepers noticed the presence of wasps at the entrance.

Hive 2

By the time hive 2 was opened, the weather had turned (only 17 degrees and stormy) and the bees were getting agitated. Queen was seen (still marked blue) and there were eggs across two frames, along with two frames of larvae and five frames of wall-to wall brood. One brood frame was solid with stores including honey and pollen. An empty frame from the back of the box was moved to the front to provide more space for the queen to lay in.

Of the two supers on hive 2, one was full and the other only recently added.

It was noted that one brood frame needed replacing due to the age of the comb and potential for disease and pathogens. This frame was changed on 9 August.

Queen is a very good layer and may be another candidate for next year’s breeding programme.

Hive 4 (position 8)

Those present at the apiary meeting on 17 July will remember this hive had recently swarmed. One frame supporting seven impressive-looking sealed cells was transferred to a nuc and the remaining more-developed cells were removed leaving only two, less developed ones in the hive.

By removing the more developed cells, the bees are given more time to settle down which reduces the urge to issue secondary (cast) swarms.


Upon opening, Denise and Richard were lucky enough to witness a ‘tooting‘ queen which is a very rare privilege indeed. Tooting is a signal emitted by recently emerged virgin queens by pressing themselves against the comb and vibrating their flight muscles without moving their wings. It starts with one or two pulses of about one second’s duration followed by a number of short pulses of about a quarter of a second.

The emerged virgin is ‘tooting’ to announce her presence in the hive and is answered by the other mature virgins still confined in their queen cells. Those still confined, use another piping sound called quacking. “When several confined queens are present in the nest, a chorus of synchronized quacking follows each tooting” (Wenner 1962; Michelsen et al. 1986).

The role of these piping signals (which are used in the period after a primary swarm has issued) is to let everybody in the hive know what’s happening! Virgin queens meeting on the comb will fight to the death (and the first to emerge is sometimes allowed to kill her sisters in their cells) but a strong colony will keep virgins apart (and delay the emergence of those trapped in their cells) so that cast swarms can issue, each led by one of the virgin queens. For more information on the piping signals click here.

The inspection of the colony revealed that only one cell had hatched. The second cell was still sealed however the queen could be seen emerging and would have got out within the hour. The frame containing this cell was moved to Hive 3 (see below) to solve a queenless situation in that hive.

Otherwise the colony had plenty of stores and capped brood awaiting hatching. The hive will be checked in 14 days to confirm a laying queen. It was also noted that the hive stand needs repairing and this might be a job for the apiary day in September.

Hive 3

This National deep belonging to Howard was also inspected at the apiary meeting. It was in a sorry story indeed! The hive was moved from Botley in an emergency last month and had been queenless for sometime. It was due to be merged with another colony this week using the newspaper method.

Upon opening, Richard and Denise noticed erratic wax building over several frames, also the bees had removed foundation from other frames for use elsewhere. They noticed all the bees were older, some with torn wings and they attributed the erratic wax building to bees that had passed the age where they can produce wax effectively. Another contributory factor could have been the foundation used was dry and past its best.

Nevertheless, the bees were busy filling the frames with honey; some of it seemed older suggesting they don’t need it currently; no brood to feed! There was also lots of stored pollen.

The supers were generally empty but there was a small amount of freshly-stored nectar.

Denise and Richard remembered the hatching queen from hive 4 and transferred her to this hive on the frame of capped brood supporting her cell. The additional brood will give the colony a population boost. They both crossed their fingers, hoping that the queen still had sufficient time before fully emerging to be accepted by her new colony. Thank you both very much for your efforts!

A note was made that a standard size frame had been added to the deep brood box which will need replacing in time. The hive will be checked again in 14 days to confirm a laying queen and several brood frames will need replacing as the bees have filled all the laying space with honey and pollen.

Hive 6

This hive was created from the frames taken from hive 4 at the apiary meeting on 17 July. There were seven sealed cells on the frame at the time and a three frame nuc was created using a small patch of brood and stores. We knew the nuc was light on stores at the time of splitting (there were no brood frames with stores and the nuc did not accommodate a super) so so Louise transferred the bees to a full size hive a week later to facilitate feeding.

Hive 6, position 10

Denise and Richard found no activity at the front of the hive which initially led them to believe the hive was empty but on opening, they found brace comb in the roof (which was removed) and a viable population.

Most of the bees would not be old enough to have commence their foraging duties and any flyers that were transferred on the 17 July would have likely returned to the parent colony.

There was no sealed brood or eggs but they saw the remnants of two (of the seven) queen cells which had been mostly dismantled. The bees were grumpy (another sign of not being queenright) and stung Richard multiple times on both hands. He was last seen running down Mayhill Road in a South-easterly direction and if anyone sees him please let us know as we’re concerned for his welfare!

The empty feeder (and eke) was removed for cleaning. The hive will be checked again in seven days for signs of a mated queen. Eggs can be added from another colony if necessary.

One thought on “Swanmore apiary report – August 2021

Leave a Reply

Fill in your details below or click an icon to log in:

WordPress.com Logo

You are commenting using your WordPress.com account. Log Out /  Change )

Twitter picture

You are commenting using your Twitter account. Log Out /  Change )

Facebook photo

You are commenting using your Facebook account. Log Out /  Change )

Connecting to %s

%d bloggers like this: